Aqueous hydrochloric acid will react with solid sodium hydroxide to produce aqueous sodium chloride and liquid water . Suppose 28.8 g of hydrochloric acid is mixed with 20. g of sodium hydroxide. Calculate the minimum mass of hydrochloric acid that could be left over by the chemical reaction. Be sure your answer has the correct number of significant digits.
1
Expert's answer
2019-09-27T04:46:08-0400
"HCl + NaOH = NaCl + H_2O""n(HCl) = n(NaOH)"

According to the data from the conditions of the assignment, we have one significant figure in the calculations.

"n(HCl) = {m(HCl) \over M(HCl)} = {28.8g \over 36.5g/mol} = 0.8 mol"

"n(NaOH) = {m(NaOH) \over M(NaOH)} = {20.0g \over 40.0g/mol} = 0.5mol"

Based on a 1: 1 stoichiometric ratio, 0.5 mol perchloric acid should be consumed in a reaction with 0.5 mol sodium hydroxide.

Then, the amount of unreacted hydrochloric acid substance:


"n(HCl)_{remainder} = 0.8mol-0.5mol = 0.3mol"

Accordingly, the minimum mass of the remaining acid:

"m(HCl) = n(HCl)*M(HCl) = 0.3mol*36.5g/mol = 10.95g \approx 11.0 g"

Answer: minimum 10.95 g (about 11.0 g)